OpenStudy (anonymous):

find critical numbers of the function f(x)=(x^-5)(lnx)

OpenStudy (freckles):

You need to use product rule to find f'(x) since you have a product for f

OpenStudy (anonymous):

so f'(x)= (x^-5)(e^x)+(lnx)(-5x^-6)

OpenStudy (freckles):

derivative of ln(x) w.r.t x is 1/x

OpenStudy (freckles):

\[f'(x)=x^{-5} \frac{1}{x}+\ln(x)(-5x^{-6}) \\ f'(x)=x^{-6}+\ln(x)(-5x^{-6})\] You can factor f'

OpenStudy (anonymous):

then i just that equal to zero to find the critical values?

OpenStudy (freckles):

yep

OpenStudy (anonymous):

thanks for the help i had a brain fart for a moment and forgot the log rules